### v1

@@ Abstract
Load balancing—the allocation of work across parallel resources to reduce delay, energy and cost—is a pervasive challenge in science and engineering, from large-scale simulation and data processing to cloud and manufacturing operations. Motivated by the emerging bottleneck in large language model (LLM) serving, we study a particularly stringent regime of load balancing that arises in barrier-synchronized, stateful systems: work cannot be freely migrated and progress is gated by the slowest participant at each step, so heterogeneity and temporal drift in workloads create persistent stragglers and substantial idle time. LLM serving under data-parallel decoding provides a prominent modern instance: in production traces, barrier-induced idle can exceed 40% of compute time per decode step. Here we develop a universal load-balancing principle, which admits a step-wise finite-horizon integer-optimization formulation and yields worst-case guarantees: across LLM decode models and a broader class of non-decreasing workload drift processes, it reduces long-run imbalance by a factor that grows with batch size and system scale. Extensive experiments corroborate the theory, showing substantial improvements in throughput and latency together with reductions in energy consumption. These results provide a general, theoretically grounded framework for load balancing, with immediate implications for sustainable LLM serving and broad relevance to other synchronization-gated resource-allocation problems.
@@ 1 Introduction
In LLM serving, inference unfolds in two stages: a prefill stage that encodes the input prompt and initializes internal state, followed by a decode stage that generates the response one token at a time. In modern deployments, decoding is executed in parallel across many computational workers, and each token step typically includes a coordinated stage across workers (e.g., collective communication in model-parallel execution), which induces barrier synchronization: the step completes when the slowest worker finishes. As active requests evolve and complete at different times, per-worker workloads drift, creating persistent stragglers and forcing faster workers to idle. Figure 1 quantifies the resulting inefficiency in a real industrial trace: both the mean and median barrier-induced idle time exceed 40% per decode step, wasting more than two-fifths of aggregate compute on average. This inefficiency scales directly into energy cost at fleet scale. Because GPUs draw substantial power even while waiting at synchronization barriers—idle time consumes energy [ozcan2025quantifying] without producing useful computation—the 40% barrier-induced idle in Figure 1 represents not merely wasted time but wasted watts. Figure 2 demonstrates this concretely: reducing decode-stage imbalance via the approach introduced in this paper yields a 3.4% reduction in total energy consumption relative to the default serving policy. At the scale of modern LLM deployments—where inference alone can consume gigawatt-hours of electricity daily—even single-digit percentage improvements translate to megawatt-hour savings with proportionate reductions in cooling load and carbon footprint. Moreover, as we establish theoretically, the improvement factor grows with system scale, implying that larger production deployments stand to benefit even more substantially. We next unpack the decode-side mechanics that create this imbalance and clarify why routing decisions are both critical and difficult.
@@
Figure 1: Workload imbalance and idle time in 436436 decoding steps over 33 minutes from a real industrial trace. (a): Dynamic total workload across 32 GPUs. At each step, the completion time for all GPUs is dictated by the heaviest workload unit, demonstrating significant and persistent workload imbalance. (b): Per-step idle time resulting from this imbalance. The yellow bars represent the average percentage of time that GPUs spend waiting during each decode step. With a mean (and median) idle time of 40% (and 41%), more than two-fifths of the aggregate computational resources are wasted on average per step.
@@
Figure 2: Instantaneous GPU power consumption over time under the default policy and the approach introduced in this paper (BF-IO). Total energy, computed as the time integral of instantaneous power, is 396 kJ for the default and 383 kJ for BF-IO —a 3.4% reduction. BF-IO draws slightly higher instantaneous power (reflecting improved GPU utilization from reduced barrier idling), but completes the same workload in less time; the resulting shorter integration window yields lower total energy consumption.
@@ 2.2 Provable theoretical guarantees under worst-case settings
Guaranteed improvement in the LLM serving model.
@@ 2.3 Resolving a decode-stage load balancing bottleneck in LLM serving
We establish that decode-stage load imbalance constitutes a dominant efficiency bottleneck in production-scale LLM serving with large-scale EP/TP synchronization. In such deployments, each token step is barrier-gated by the slowest data-parallel worker; heterogeneity in resident KV footprints therefore translates directly into idle time on faster workers while they await synchronization. Our analysis of real industrial traces quantifies this inefficiency: Figure 1(b) shows that the average fraction of compute time lost per decode step exceeds 40%, representing more than two-fifths of aggregate GPU resources wasted at barriers, which results in a 3.4% energy waste. Crucially, the magnitude of this waste varies sharply with arrival patterns and request characteristics, indicating that improvements tied to any single trace or dataset are insufficient—a finding that motivates our focus on a universal principle with worst-case guarantees.
@@
We evaluate BF-IO through extensive numerical simulations to assess whether the theoretical improvements translate into practical efficiency gains. We evaluate BF-IO on (i) public datasets and (ii) proprietary industrial workloads under matched serving configurations (Section 6). Across these settings, BF-IO consistently reduces imbalance and improves system efficiency relative to the default baselines used in real industries. Concretely, on the public workload, BF-IO improves throughput by 14%, reduces latency by 13%, and reduces energy consumption by 3.4%.
@@ 4 Method: A Universal Load Balance Principle
Figure 4: Intuition for selecting the window length $H$, evaluated across values from 0 to $500$. This analysis uses an implementation of the Algorithm BF-IO in the LLM serving context. Further implementation and experimental details are provided in Sections 6.
@@ 5.1 Theoretical Guarantee of BF-IO in General Cases
$\mathbf{IIR}\;=\;\inf_{\mathcal{I}}\;\frac{\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})}{\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})}\;\geq\;c\,\frac{p\,\sigma_{s}}{s_{\max}}\,\frac{G}{G-1}\,\sqrt{B\log G}\;=\;\Omega\!\big(\sqrt{B\log G}\big).$ (5)
@@ 6 Experiments
We implement a numerical simulation framework to evaluate different scheduling strategies for GPU-based request processing. This section describes the experimental setup, evaluation metrics and numerical result of different routing strategies.
6.1 Request Generation
To simulate realistic LLM inference workloads, we use request traces from the BurstGPT dataset [wang2025burstgpt]. Each request is characterized by a triplet $(s_{i},o_{i})$, where $s_{i}$ denotes the prompt (input) length, $o_{i}$ the actual output length.
Dataset.
The BurstGPT dataset captures production-level LLM serving patterns, including the heavy-tailed distribution of output lengths commonly observed in real deployments. We extract the input_tokens and output_tokens fields from the dataset to construct our simulation workload. This approach ensures that our experiments reflect realistic request characteristics, such as the variability and correlation structure present in actual user queries.

### v2

@@ Abstract
Over 40% of computational power in Large Language Model (LLM) serving systems can be systematically wasted—not from hardware limits, but from load imbalance in barrier-synchronized parallel processing. When progress is gated by the slowest worker at each step, heterogeneous and evolving workloads create persistent stragglers; faster workers idle while drawing power, producing nothing. In large language model inference alone, this translates to gigawatt-hours of wasted electricity daily. Here we develop a universal load-balancing principle for barrier-synchronized systems with non-migratable state. We prove worst-case theoretical guarantees: imbalance reduction grows with system scale, and the resulting energy savings can exceed 52% for modern hardware at fleet scale. Experiments corroborate the theory, demonstrating 28% energy reduction alongside substantial throughput and latency improvements. Formulated as an online integer optimization with provable guarantees, the principle extends beyond LLM serving to broad classes of barrier-synchronized parallel systems, establishing a theoretical foundation for sustainable high-performance computing.
@@ 1 Introduction
In LLM serving, inference unfolds in two stages: a prefill stage that encodes the input prompt and initializes internal state, followed by a decode stage that generates the response one token at a time. In modern deployments, decoding is executed in parallel across many computational workers, and each token step typically includes a coordinated stage across workers (e.g., collective communication in model-parallel execution), which induces barrier synchronization: the step completes when the slowest worker finishes. As active requests evolve and complete at different times, per-worker workloads drift, creating persistent stragglers and forcing faster workers to idle. Figure 1 quantifies the resulting inefficiency in a real industrial trace: both the mean and median barrier-induced idle time exceed 40% per decode step, wasting more than two-fifths of aggregate compute on average. This inefficiency scales directly into energy cost at fleet scale. Because GPUs draw substantial power even while waiting at synchronization barriers—idle time consumes energy [ozcan2025quantifying] without producing useful computation—the 40% barrier-induced idle in Figure 1 represents not merely wasted time but wasted watts. Figure 2 demonstrates this concretely: reducing decode-stage imbalance via the approach introduced in this paper yields a 28.2% reduction in total energy consumption relative to the default serving policy over this trace, with the improvement growing as system scale increases. At the scale of modern LLM deployments—where inference alone can consume gigawatt-hours of electricity daily—such reductions translate to megawatt-hour savings with proportionate reductions in cooling load and carbon footprint. We next unpack the decode-side mechanics that create this imbalance and clarify why routing decisions are both critical and difficult.
@@
Figure 1: Workload imbalance and idle time in 436436 decoding steps over 33 minutes from a real industrial trace. (Left): Dynamic total workload across 32 GPUs. At each step, the completion time for all GPUs is dictated by the heaviest workload unit, demonstrating significant and persistent workload imbalance. (Right): Per-step idle time resulting from this imbalance. The yellow bars represent the average percentage of time that GPUs spend waiting during each decode step. With a mean (and median) idle time of 40% (and 41%), more than two-fifths of the aggregate computational resources are wasted on average per step.
@@
Figure 2: Comparison of energy and power consumption over the default policy and the approach introduced in this paper (BF-IO). (Left): Instantaneous GPU power consumption over time under the default policy (orange curve) and BF-IO (blue curve). Total energy, computed as the time integral of instantaneous power, is 29.1 MJ for the default and 20.9 MJ for BF-IO —a 28.2% reduction. BF-IO draws slightly higher instantaneous power (reflecting improved GPU utilization from reduced barrier idling), but completes the same workload in less time; the resulting shorter integration window yields lower total energy consumption. (Right): Energy consumption over increasing system scale (number of GPUs). The energy reduction percentage keeps increasing when the system becomes larger.
@@ 2.2 Provable theoretical guarantees under worst-case settings
Guaranteed load balancing improvement in the LLM serving model.
@@
From imbalance reduction to energy savings.
Beyond efficiency metrics, we establish a rigorous theoretical connection between load balancing and energy consumption. Theorem 4 proves that any guaranteed improvement in imbalance translates into an explicit guaranteed improvement in total synchronized-phase energy, with constants depending on GPU power model parameters and request distribution. Instantiating this result with our imbalance guarantees for BF-IO yields Corollary 1: as system scale $G\to\infty$, the energy saving fraction approaches a hardware-dependent constant that exceeds 52% for modern GPUs. This provides, to our knowledge, the first worst-case theoretical guarantee linking scheduling policy to energy efficiency in LLM serving, establishing a principled foundation for sustainable inference at scale.
@@ 2.3 Resolving a decode-stage load balancing bottleneck in LLM serving
We establish that decode-stage load imbalance constitutes a dominant efficiency bottleneck in production-scale LLM serving with large-scale EP/TP synchronization. In such deployments, each token step is barrier-gated by the slowest data-parallel worker; heterogeneity in resident KV footprints therefore translates directly into idle time on faster workers while they await synchronization. Our analysis of real industrial traces quantifies this inefficiency: Figure 1 shows that the average fraction of compute time lost per decode step exceeds 40%, representing more than two-fifths of aggregate GPU resources wasted at barriers, which results in a 28.2% energy waste. Crucially, the magnitude of this waste varies sharply with arrival patterns and request characteristics, indicating that improvements tied to any single trace or dataset are insufficient—a finding that motivates our focus on a universal principle with worst-case guarantees.
@@
We evaluate BF-IO through extensive numerical simulations to assess whether the theoretical improvements translate into practical efficiency gains. We evaluate BF-IO on (i) public datasets and (ii) proprietary industrial workloads under matched serving configurations (Section 6). Across these settings, BF-IO consistently reduces imbalance and improves system efficiency relative to the default baselines used in real industries. Concretely, on the public workload, BF-IO improves throughput by 90%, reduces latency by 44%, and reduces energy consumption by 28%. The proprietary workload has been verified by a major LLM developer industry in China.
@@ 4 Method: A Universal Load Balance Principle
Figure 4: Intuition for selecting the window length $H$, evaluated across values from 0 to $100$. This analysis uses an implementation of the Algorithm BF-IO in the LLM serving context. Further implementation and experimental details are provided in Sections 6.
@@ 5.1 Theoretical Guarantee of BF-IO in General Cases
$\mathbf{IIR}\;=\;\inf_{I\in\mathcal{I}}\left(\lim_{K\to\infty}\frac{\mathbb{E}\bigl[\text{AvgImbalance}(\textsf{FCFS};I)\bigr]}{\mathbb{E}\bigl[\text{AvgImbalance}(\textsf{BF-IO};I)\bigr]}\right)\;\geq\;c\,\frac{p\,\sigma_{s}}{s_{\max}}\,\frac{G}{G-1}\,\sqrt{B\log G}\;=\;\Omega\!\big(\sqrt{B\log G}\big).$ (5)
@@
5.2 Theoretical Guarantee of BF-IO on Energy Consumption
$E\;=\;\int_{0}^{T}P\bigl(\mathrm{mfu}(t)\bigr)\,dt,$ (6)
where $\mathrm{mfu}(t)$ denotes the Model FLOPs Utilization (mfu), i.e., the ratio of observed computational throughput to the GPU’s theoretical peak FLOPs capacity. Empirical measurements in [ozcan2025quantifying] validate that instantaneous GPU power is a sublinear function of utilization:
$P(\mathrm{mfu})\;=\;P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\left(\frac{\mathrm{mfu}}{\mathrm{mfu}_{\mathrm{sat}}}\right)^{\gamma},\qquad\gamma\in(0,1),$ (7)
where $P_{\mathrm{idle}}$ is idle power, $P_{\mathrm{max}}$ is peak power, $\mathrm{mfu}_{\mathrm{sat}}$ is a saturation threshold, and $\gamma<1$ captures the inefficiency that GPUs draw substantial power even at low utilization.
Synchronized phase and utilization model.
To isolate the impact of barrier-induced waiting, we focus on the synchronized phase of each decoding step (e.g., the attention computation) during which all workers wait for the slowest one. Other phases (e.g., all-to-all EP/TP communication) can be incorporated additively but are omitted here for clarity.
Recall from Section 3 that at discrete step $k$ the instantaneous workload on worker $g$ under policy $\pi$ is $L_{g}(k)$, and $L_{g^{*}}(k)=\max_{g\in[G]}L_{g}(k)$. The wall-clock duration of the synchronized attention phase in step $k$ is linear in the maximal workload, namely, $\tau_{k}(\pi)\;=\;\kappa_{\mathrm{ATT}}\cdot L_{g^{*}}(k)$, for some constant $\kappa_{\mathrm{ATT}}>0$ (seconds per unit workload). During this synchronized phase, worker $g$ performs useful attention computation for $\kappa_{\mathrm{ATT}}L_{g}(k)$ seconds and waits at the barrier for $\kappa_{\mathrm{ATT}}(L_{g^{*}}(k)-L_{g}(k))$ seconds. Hence its utilization fraction within the synchronized phase is
$u_{g}(k)\;:=\;\frac{L_{g}(k)}{L_{g^{*}}(k)}\in[0,1].$ (8)
As the utilization fraction equals to the instant throughput fraction, $\frac{\mathrm{mfu}_{g}(t)}{\mathrm{mfu}_{\mathrm{sat}}}$, we have
$\frac{\mathrm{mfu}_{g}(t)}{\mathrm{mfu}_{\mathrm{sat}}}\;=\;u_{g}(k),\qquad t\ \text{during the synchronized phase of step }k.$ (9)
Under (7)–(9), the per-worker power in step $k$ equals $P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}.$
Total synchronized-phase energy.
Fix an arrival instance $\mathcal{I}$ and a policy $\pi$. Let $K_{\pi}(\mathcal{I})$ denote the number of decoding steps until all requests in $\mathcal{I}$ complete under $\pi$. The total energy consumed in the synchronized attention phase across all $G$ workers is
$E(\pi;\mathcal{I})$ $:=\sum_{g=1}^{G}\int_{0}^{T_{\pi}}P\!\left(\mathrm{mfu}_{g}(t)\right)\,dt$
$=\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}\tau_{k}(\pi)\Bigl(P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}\Bigr)$
$=\kappa_{\mathrm{ATT}}\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}L_{g^{*}}(k)\sum_{g=1}^{G}\Bigl(P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}\Bigr),$ (10)
where $T_{\pi}=\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\tau_{k}(\pi)$ is the corresponding makespan of the synchronized phase.
Define the policy-independent total attention workload of the instance $\mathcal{I}$ as:
$W(\mathcal{I})\;:=\;\sum_{i\in\mathcal{I}}\sum_{j=1}^{o_{i}}w_{i}^{(j)}\;=\;\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}L_{g}(k),$ (11)
where the second equality holds for every policy $\pi$ because each request contributes its full workload profile exactly once regardless of assignment timing or placement. Also define the cumulative imbalance up to completion:
$\mathrm{ImbTot}(\pi;\mathcal{I})\;:=\;\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\mathrm{Imbalance}(k;\pi)\;=\;\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}\bigl(L_{g^{*}}(k)-L_{g}(k)\bigr).$ (12)
Finally, for a baseline policy $\pi_{0}$, define the normalized imbalance level
$\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})\;:=\;\frac{\mathbb{E}\bigl[\mathrm{ImbTot}(\pi_{0};\mathcal{I})\bigr]}{\mathbb{E}\bigl[W(\mathcal{I})\bigr]}.$ (13)
Energy savings from imbalance reduction.
We now show that any guaranteed improvement in imbalance yields an explicit guaranteed improvement in total synchronized-phase energy (as a percentage), with constants depending only on the power model parameters.
Theorem 4 (Energy saving guarantee from imbalance reduction).
Let $\pi_{0}$ be any baseline policy and $\pi_{1}$ any improved policy. Suppose the imbalance improvement factor satisfies
$\mathbf{IIR}\;=\;\inf_{I\in\mathcal{I}}\left(\lim_{K\to\infty}\frac{\mathbb{E}\bigl[\text{AvgImbalance}(\pi_{0};I)\bigr]}{\mathbb{E}\bigl[\text{AvgImbalance}(\pi_{1};I)\bigr]}\right)\ \geq\ \alpha\qquad\text{for some }\alpha>1.$ (14)
Define
$C_{\gamma}\;:=\;(1-\gamma)P_{\mathrm{max}}+\gamma P_{\mathrm{idle}},\qquad D_{\gamma}\;:=\;(1-\gamma)(P_{\mathrm{max}}-P_{\mathrm{idle}}),$ (15)
then the synchronized-phase energy saving fraction satisfies
$\frac{\mathbb{E}\bigl[E(\pi_{0};\mathcal{I})\bigr]-\mathbb{E}\bigl[E(\pi_{1};\mathcal{I})\bigr]}{\mathbb{E}\bigl[E(\pi_{0};\mathcal{I})\bigr]}\ \geq\ \frac{1}{P_{\mathrm{max}}/\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})+C_{\gamma}}\left(P_{\mathrm{idle}}\Bigl(1-\frac{1}{\alpha}\Bigr)\;-\;\frac{D_{\gamma}}{\alpha}\right).$ (16)
Proof sketch.
Equation (10) expresses $E(\pi;\mathcal{I})$ as a sum over steps of the barrier duration $\kappa_{\mathrm{ATT}}L_{g^{*}}(k)$ times power. Writing $u_{g}(k)^{\gamma}=u_{g}(k)+(u_{g}(k)^{\gamma}-u_{g}(k))$ and using $L_{g^{*}}(k)u_{g}(k)=L_{g}(k)$ yields an exact decomposition:
$E(\pi;\mathcal{I})=\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I})+\text{(concavity correction)}.$
Since $\gamma\in(0,1)$, $u^{\gamma}$ is concave on $[0,1]$; the tangent bound at $u=1$ implies that the concavity correction is nonnegative and at most $\kappa_{\mathrm{ATT}}D_{\gamma}\,\mathrm{ImbTot}(\pi;\mathcal{I})$. This shows $E(\pi;\mathcal{I})$ is sandwiched between two affine functions of $\mathrm{ImbTot}(\pi;\mathcal{I})$ with the same policy-independent baseline term $\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})$. Combining these bounds with the imbalance ratio assumption (14) and normalizing by $W(\mathcal{I})$ yields (16). The complete proof can be found in Appendix C.4.
Remark 1 (Instantiation to $\pi_{0}=\textsf{FCFS}$ and $\pi_{1}=\textsf{BF-IO}$ in our LLM model).
In the inhomogeneous LLM decode model of Section 5, if we take $\pi_{0}=\textsf{FCFS}$ and $\pi_{1}=\textsf{BF-IO}$, then (14) holds with $\alpha=c\,\frac{p}{s_{\max}}\,\sqrt{\sigma_{s}^{2}+\frac{1-p}{p^{2}}}\;\frac{G}{G-1}\,\sqrt{B\log G}$, and thus Theorem 4 yields a nontrivial constant lower bound on the synchronized-phase energy saving fraction once $\mathbf{IIR}$ is sufficiently large. Moreover, under the overloaded regime and geometric output lengths, the normalized imbalance level $\eta_{\mathrm{sum}}(\textsf{FCFS};\mathcal{I})$ admits an explicit lower bound:
$\eta_{\mathrm{sum}}(\textsf{FCFS};\mathcal{I})\ \gtrsim\ \frac{\sigma_{\mathrm{snap}}}{\mu_{s}+\frac{1-p}{p}}\sqrt{\frac{\log G}{B}},\qquad\sigma_{\mathrm{snap}}^{2}=\sigma_{s}^{2}+\frac{1-p}{p^{2}}.$ (17)
Plugging (17) and $\alpha=c\,\frac{p}{s_{\max}}\,\sqrt{\sigma_{s}^{2}+\frac{1-p}{p^{2}}}\;\frac{G}{G-1}\,\sqrt{B\log G}$ into (16) yields an explicit energy saving guarantee in terms of $(p,\sigma_{s},s_{\max},B,G)$ and $(P_{\mathrm{idle}},P_{\mathrm{max}},\gamma)$. The complete proof of Inequality (17) can be found in Appendix C.5
Corollary 1 (Asymptotic energy saving at scale).
Under the conditions of Remark 1, as the number of workers $G\to\infty$, the guaranteed energy saving fraction of BF-IO relative to FCFS converges to
$\lim_{G\to\infty}\frac{\mathbb{E}\bigl[E(\textsf{FCFS};\mathcal{I})\bigr]-\mathbb{E}\bigl[E(\textsf{BF-IO};\mathcal{I})\bigr]}{\mathbb{E}\bigl[E(\textsf{FCFS};\mathcal{I})\bigr]}\ \geq\ \frac{P_{\mathrm{idle}}}{(1-\gamma)P_{\mathrm{max}}+\gamma P_{\mathrm{idle}}}.$ (18)
Remark 2 (Concrete instantiation for modern GPUs).
For NVIDIA A100 GPUs, empirical measurements in [ozcan2025quantifying] report $P_{\mathrm{idle}}=100$W, $P_{\mathrm{max}}=400$W, and $\gamma=0.7$. Substituting into (18) in Corollary 1:
$\frac{P_{\mathrm{idle}}}{(1-\gamma)P_{\mathrm{max}}+\gamma P_{\mathrm{idle}}}\;=\;\frac{100}{0.3\times 400+0.7\times 100}\;=\;\frac{100}{190}\;\approx\;52.6\%.$
That is, in large-scale deployments, the principled load balancing introduced in this work can reduce synchronized-phase energy consumption by more than half compared to the widely-deployed FCFS baseline. This asymptotic guarantee underscores the sustainability impact of addressing the decode-stage load imbalance bottleneck at fleet scale.
@@ 6 Experiments
We evaluate the proposed routing strategies through discrete-event simulation of GPU-based LLM request processing. This section describes the experimental setup, evaluation metrics, and numerical results.
6.1 Dataset and Request Generation
To simulate realistic LLM inference workloads, we use request traces derived from LongBench [bai2024longbench], a challenging benchmark designed for long-context language models. LongBench encompasses diverse long-text application scenarios—single-document and multi-document question answering, summarization, few-shot learning, synthetic tasks, and code completion—and reflects production-relevant settings in which input prompts are substantially longer than typical benchmarks. Each request is characterized by a pair $(s_{i},o_{i})$, where $s_{i}$ denotes the prefill (input prompt) length and $o_{i}$ the decode (output) length; the resulting distributions are shown in Figure 6. Requests arrive according to a stationary Poisson process at a rate exceeding the system’s processing capacity, ensuring that the system operates in the overloaded regime central to our theoretical analysis. For comparison, we also evaluate on BurstGPT [wang2025burstgpt], a lighter-load trace; these results appear in Appendix D.2.
Figure 6: Prefill and decode length distribution of LongBench.